OpenStudy (anonymous):

You toss 5 coins. What is the probability that all 5 show heads?

OpenStudy (anonymous):

1 32 1 16 1 12 1 4

OpenStudy (mathstudent55):

The 5 coin tosses are independent events. The outcome of each event does not influence any other events. Find the probability of each event and multiply them all together.

OpenStudy (mathstudent55):

Also, you can type fractions this way: A. 1/32 B. 1/16 C. 1/12 D. 1/4

OpenStudy (anonymous):

so whats the answer ? because i don't get it

OpenStudy (mathstudent55):

When you toss one coin, there are 2 possible outcomes. You are interested in one outcome. The probability of an outcome is: number of desired outcomes/total number of outcomes If you want a heads, your number of desired outcomes is 1 Since a coin can fall as heads or tails, there are two possible outcomes. The probability of getting heads is 1/2 which is the 1desired outcome divided by the 2 possible outcomes. Ok so far?

OpenStudy (anonymous):

0.25 ?

Mehek (mehek14):

each coin has \(\dfrac{1}{2}\) of getting heads flipping 5 coins has a chance of \(\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}\)

OpenStudy (mathstudent55):

No. For each toss, the probability of heads is 1/2 Since all tosses are independent events, you multiply all their probabilities together: 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 1/32